OpenStudy (anonymous):

Just need someone to check my answers. find f ^-1(x) f(x)= 3squareroot x-3 = f^-1(x)= x^3+3 Simply the expression, if possible log 0.0001 = -4 A theater has 29 seats in the first row, 42 seats in the second row, 45 seats in the third row, and so on. a. Can the number of seats in each row by modeled by an arithmetic or geometric sequence? b. Write the general terms for a sequence a n that gives the number of seats in row n. = n19=29+(19-1)3 c. How many seats are there in row 19 = 83

OpenStudy (anonymous):

@satellite73 or @amistre64

OpenStudy (anonymous):

1) \[y=3\sqrt{x+3}\implies{y\over3}=\sqrt{x+3}\implies{y^2\over9}=x+3\\\implies x=\boxed{\Large{y^2\over9}-3=f^{-1}(x)}\]

OpenStudy (anonymous):

is my understanding of your f(x) is correct, then your inverse for that is wrong.

OpenStudy (anonymous):

ooh you had it this way? 1) \[y=\sqrt[3]{x-3}\implies y^3=x-3\implies x=\boxed{y^3+3=f^{-1}(x)}\] 2) correct 3) you mean "55" in the third row? then yes.

OpenStudy (anonymous):

How did you get 55?

OpenStudy (anonymous):

in the third question, you used arithmatic progression. I do not see it in the way you wrote.